solve the linear system: y=2x+2 equ 1 x+3y=41 equ2
Answer by mananth(16946)

y=2x+2
y-2x=2..........1
x+3y=41.........2
multiply (1)by 1
Multiply (2) by 2
-2x+y=2
2x+6y=82

Add the two equations

7y=84

/7

y=12
Plug value of y in
-2x+y=2
-2x+12 =2
-2x=-10
/-2
x=5